could you help me with this equation? it is from a question we had in class.
The focus of (x-1)^2 = (y-2) is_____.
i said the answer is (-2,-1)
can you tell me if i got it right? thanks for any help.

Answer by Nate(3500)

(x - 1)^2 = y - 2
(x - 1)^2 + 2 = y
v(1,2) and a = 1
p = 1/(4a) = 1/4
(1,2.25)